| 12/21/2004 | |||
| 12:22 | Data connected | KC | |
| 12:32 | SCI_SW_HV 0x04, 1250 | KC | |
| 12:56 | SCI_SW_HV 0x04, 1500 | HP | |
| 1:14 | SCI_SW_HV 0x04, 1750 | KC | |
| 1:22 | SCI_SW_HV 0x04, 2000 | KC | |
| 1:25 | We reached 6.09kv for MB5 of Cal. Successfully and no trip at all so far!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! | KC | |
| 3:15 | Data stopped for 10 minutes due to window time | KC | |
| 4:30 | Data will be stopped for 30 minutes due to window time | KC | |
| 6:40 | Cal trigger thresholds to 82 DAC | KC | |
| 6:46 | Playback started | KC | |
| 7:47 | ESC said will not get data in 30 min | HQ | |
| 8:16 | Start getting data | HQ | |
| 8:26 | Cal trigger thresholds to 80 DAC | TG | |
| 11:45 | Cal trigger thresholds to 75 DAC | OG | |
| 11:45 | Ramping hodo MBs 2 & 5 to 9 kV | OG | |
| 11:51 | MB5 Bias and HV tripped | TG | |
| 11:58 | go to stanby mode, and daq restart | TG | |
| 11:58 | Hodo MB03 tripped incl bias and power | OG | |
| 12:01 | Forward link lost for next 50 minutes, try then | OG | |
| to recover power on hodo MB03 | |||
| Hodo MB05 now 7.5kV, will ramp to 9kV later | |||
| 12:50 | Hodo MB5 turn off and turned on | ||
| 13:03 | ramping HOD MB05 HV to 2750 | HQ | |
| CAL MB05 bias off then bias on | |||
| ramping CAL MB05 HV to 750 (Void) | |||
| 13:15 | SCI_SW_HV_LEVEL 0x0F, 2250 (MB8 HDS) | KC | |
| SCI_SW_HV_LEVEL 0x0C, 3000 (MB5 HDS) | |||
| HOD MB08 HV set to 7.5KV | |||
| 13:41 | SCI_SW_HV_LEVEL 0x0F, 2750 (MB8) | KC | |
| 14:00 | SCI_SW_HV_LEVEL 0x0F, 3000 | ||
| Tripped MB8 a few minutes after 3000 | |||
| wait for the next command window | |||
| 14:55 | command window open | KC | |
| Power off and on for MB8 of HDS | |||
| 15:00 | window closed | KC | |
| 15:15 | window open | KC | |
| Turned off and on bias of MB8 of HDS | |||
| Turned on HV of MB8 of HDS | |||
| Started ramping up to 250 of MB8 of HDS | |||
| 15:28 | SCI_SW_HV_LEVEL 0x0F, 500 | KC | |
| 15:33 | SCI_SW_HV_LEVEL 0x0F, 750 | KC | |
| 15:37 | SCI_SW_HV_LEVEL 0x0F, 1000 | HQ | |
| 15:42 | SCI_SW_HV_LEVEL 0x0F, 1250 | ||
| 17:06 | Sonny restarted commanding
programs on CDPS1 after changing time zone to GMT. |
||
| 17:33 | SCI_SW_HV_LEVEL 0x0F, 1500 | KC | |
| 17:35 | cal_spars_update.stol | KC | |
| Cal trigger thresholds to 72 DAC | |||
| 17:40 | SCI_SW_HV_LEVEL 0x0F, 1750 | KC | |
| 17:48 | SCI_SW_HV_LEVEL 0x0F, 2000 | ||
| 17:52 | SCI_SW_HV_LEVEL 0x0F, 2250 | ||
| 17:57 | SCI_SW_HV_LEVEL 0x0F, 2500 | ||
| 18:36 | Window open | KC | |
| SCI_SW_HV_LEVEL 0x0F, 2750 | |||
| Turn off and on Cal trigger Bias - now OK | |||
| Cal trigger threshold kept 72 because | |||
| the rate changing 1.0 to 3.0 and data rate is | |||
| 25 to 40. And sometime buffer showing | |||
| 0.2 % building. | |||
| 18:43 | SCD ch#2880 threshold level changed from 40000 to 70000 | JH | |
| play-back data 18:50~18:54 | |||
| 19:00 | connection of data flow is stopped | ||
| 19:15 | reconnection | ||
| 19:49 | SCD ch#2880 threshold level changed from 40000 to 70000 | JH | |
| command went through, but it's not changed->why? | |||
| 19:50 | HDS MB08 is tripped | JH | |
| HV->0, power off & power on | |||
| 19:56 | SCI_SW_OPMODE 0 | JH | |
| SCI_SNIO_CONTROL 21,0 | |||
| Execute script daq_restart.stol | |||
| 20:06 | HDS MB08 bias on | JH | |
| 20:09 | HDS MB08 HV on | ||
| 20:10 | SCI_HDS_HV_LEVEL 0x14, 1000 | ||
| 20:19 | HDS MB08 HV off | ||
| 20:20 | HDS MB08 HV on, HV level: 1000 | ||
| 20:25 | SCI_SW_HV ox14, 1000 | ||
| high voltage of HDS HB08 didn't go up-> why? | |||
| 20:26 | SCI_SW_OPMODE 0 | JH | |
| SCI_SNIO_CONTROL 21,0 | |||
| Execute script daq_restart.stol | |||
| 20:37 | SCI_SW_OPMODE 0 | SO | |
| SCI_SNIO_CONTROL 21,0 | |||
| Execute script daq_restart.stol | |||
| 21:06 | play back data is coming | JH | |
| 21:19 | SCI_SW_OPMODE 0 | JH | |
| SCI_SNIO_CONTROL 21,0 | |||
| Execute script daq_restart.stol | |||
| 21:20 | SCI_SW_HV ox14, 1000 | ||
| 21:30 | SCI_SW_OPMODE 0 | SO | |
| SCI_SNIO_CONTROL 21,0 | |||
| Execute script daq_restart.stol | |||
| 21:39 | SCI_SW_OPMODE 0 | SO | |
| SCI_SNIO_CONTROL 21,0 | |||
| Execute script daq_restart.stol | |||
| 22:23 | TDRSS quality wasn't good, so we're just getting ERROR message | JH | |
| 22:31 | The
signal was so weak not to get any data, so we just waited for ESC`s response |
SO | |
| 2:53 | |||